2x+3y=26
1x+8y=14

Answer by mananth(16946)

2 x + 3 y = 26 .............1
1 x + 8 y = 14 .............2
Eliminate y
multiply (1)by -8
Multiply (2) by 3
-16 x -24 y = -208
3 x 24 y = 42
Add the two equations
-13 x = -166
/ -13
x = 12.77
plug value of x in (1)
2 x + 3 y = 26
25.54 + 3 y = 26
3 y = 26 -25.54
3 y = 0.46
y = 0.15